2( x+3) = 10 4x - 2x + 3 = 2 x + 2 x - x

Answer by checkley71(8403)

2(x+3)=10 or 2x+6=10 or 2x=10-6 or 2x=4 or x=4/2 or x=2 answer.
-------------------------------------------------------------
4x-2x+3=2x+2x-x
2x+3=3x
2x-3x=-3
-x=-3
x=-3/-1
x=3 answer.